OpenStudy (anonymous):

Which is the exact area of the circle? A circle is shown with its radius marked 9 feet eighty-one square feet eighty-one times pi square feet eighty-one times pi squared square feet eighty-one squared times pi squared square feet

OpenStudy (tanner23456):

A = pi *r^2 A = (3.14)(9)^2 A = 81*pi ft^2 so b is your answer